OpenStudy (anonymous):

Solve this system of equations by substitution. Determine whether the system is consistent or inconsistent. 2x+y=9 y=5x+2

OpenStudy (ajprincess):

Put 5x+2 in place of y in 2x+y=9 and solve for x. Can u do that @Educate_Me?

OpenStudy (anonymous):

Yes, then I did 2(5x+2)=9 10x+4=9 -4 -4 10x=5 divide both sides by 10 , is that right ?

OpenStudy (ajprincess):

hmm no. 2x+y=9 2x+5x+2=9 7x+2=9 Nw try to solve

OpenStudy (anonymous):

I got it! I substituted the 1 I got for x into the y and got the ordered pair (1,7)

OpenStudy (ajprincess):

Perfect:)

OpenStudy (anonymous):

now this one , I am very confused.. 1/2x + 3/2y=-7 1/3y=2x-10

OpenStudy (ajprincess):

\[\frac{ 1 }{ 2}x+\frac{ 3 }{ 2 }y=-7\] \[\frac{ 1 }{ 3 }y=2x-10\]

OpenStudy (ajprincess):

R those the eqns?

OpenStudy (anonymous):

Yes.

OpenStudy (ajprincess):

what do u get when u multiply ur second equation by 3?

OpenStudy (anonymous):

umm , 3y=2x-10?

OpenStudy (ajprincess):

hmm no.